OpenStudy (anonymous):

Find the quotient. (x^2 - 4xy) ÷ x

OpenStudy (solomonzelman):

At first, answer 2 questions x² ÷ x = ? 4xy ÷ x = ?

OpenStudy (anonymous):

idk

OpenStudy (phi):

what is \[ \frac{x}{x} \]? anything divided by itself is ? what is \[ \frac{x}{x}\cdot x \] which can also be written \[ \frac{x\cdot x}{x} \] or \[ \frac{x^2}{x} \]